Hi! I need help with this Chemistry problem please:

Consider the following reaction:

A(g)⇌2B(g)


Find the equilibrium partial pressures of A and B for each of the following different values of Kp. Assume that the initial partial pressure of B in each case is 1.0 atm and that the initial partial pressure of A is 0.0 atm. Make any appropriate simplifying assumptions.

part a.)

Kp= 1.0

Enter your answers numerically separated by a comma. Express your answers using two significant figures.

PA

P A

, PB =

        

atm

part b)

Kp= 1.2×10−4

Enter your answers numerically separated by a comma. Express your answers using two significant figures.

PA

P A

, PB =

         

atm

part c)

Kp= 1.4×105

Enter your answers numerically separated by a comma. Express your answers using two significant figures.

PA

P A

, PB =

           

atm
Answer #1

                              A(g) 2B(g)              Total pressure (atm)      Qp

Initially Pressure     0             1                            1                            pB2 / pA = , Qp > kp => backward reaction

At equb.    p    1-2p    1-2p+p = 1-p

kp = pB2 / pA

a) 1 = [p/(1-p)] / [(1-2p)/(1-p)]2

=> 1 = p(1-p)/(1-2p)2 =>   4p2 - 4p +1 = p - p2       => 5p2 - 5p +1 = 0

=> p =    0.72 or 0.28

=> if p = 0.72,    At equb : P(A) = 0.72 atm , P(B) = -0.44 => not acceptable solution

& if p = 0.28 At equb: P(A) = 0.28 atm , P(B) = 0.44 atm

Similar method will be followed by next two cases.